OpenStudy (anonymous):

Either (A) an arithmetic sequence its terms do not match and n belongs to N * 1) Prove that (1/U1 * U2) + 1 (U2 * U3) + .... +1 / (A * A +1) = n / (U1 * A +1)

OpenStudy (unklerhaukus):

What is N* ?

ganeshie8 (ganeshie8):

naturals i guess \(n \in \mathbb{N}\)

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

i dont really understand the first sentence

OpenStudy (anonymous):

can you speak or understand french

OpenStudy (unklerhaukus):

i cannot speak french sorry

OpenStudy (anonymous):

ok the first sentence: (Un) an arithmetic sequence

ganeshie8 (ganeshie8):

Oh, so you mean below is the arithmetic sequence ? U1, U2, U3 ....

OpenStudy (anonymous):

yes exactly !

ganeshie8 (ganeshie8):

given that, u want to prove below ? \(\large \frac{1}{U1 * U2} + \frac{1}{U2 * U3} + .... + \frac{1 }{A * A +1} = \frac{n }{ U1 * A +1}\)

ganeshie8 (ganeshie8):

n = number of terms on left series A = ?

OpenStudy (anonymous):

yes but not A it's (Un) okéy ?

ganeshie8 (ganeshie8):

now it makes some sense. so u wanto prove below :- \(\large \frac{1}{U1 * U2} + \frac{1}{U2 * U3} + .... + \frac{1 }{U_n * U_{n+1}} = \frac{n }{ U1 * U_{n+1}}\) ?

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

So What should I do

ganeshie8 (ganeshie8):

can we use induction ?

OpenStudy (anonymous):

YeS we can

OpenStudy (anonymous):

have you found the solution

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

To prove :- \(\large \frac{1}{U1 * U2} + \frac{1}{U2 * U3} + .... + \frac{1 }{U_n * U_{n+1}} = \frac{n }{ U1 * U_{n+1}} \) rewriting the given statement :- To prove :- \(\large \frac{1}{a(a+d)} + \frac{1}{(a+d)(a+2d)} + .... + \frac{1 }{(a+(n-1)d) * (a+nd)} = \frac{n }{ a(a+nd)} \) clearly, for \(n = 1\), this is true. so, \( n=1\) is in the solution set assume the given statement is true for \(n = k\) \(\large \frac{1}{a(a+d)} + \frac{1}{(a+d)(a+2d)} + .... + \frac{1 }{(a+(k-1)d) * (a+kd)} = \frac{k }{ a(a+kd)} \)

OpenStudy (anonymous):

Did you find it

ganeshie8 (ganeshie8):

we need to the formula works for \(n = k+1\)

ganeshie8 (ganeshie8):

sum of \(k+1\) terms :- \(\large \frac{1}{a(a+d)} + \frac{1}{(a+d)(a+2d)} + .... + \frac{1 }{(a+(k-1)d) * (a+kd)} +\frac{1 }{(a+kd) * (a+(k+1)d)} \) from the induction hypothesis, sum till k terms equals k/(a(a+kd)) \(\large \frac{k }{a * (a+kd)} +\frac{1 }{(a+kd) * (a+(k+1)d)} \) \(\large \frac{k* (a+(k+1)d) }{a * (a+kd)* (a+(k+1)d)} +\frac{1*a }{a*(a+kd) * (a+(k+1)d)} \) \(\large \frac{k* (a+(k+1)d) + a }{a * (a+kd)* (a+(k+1)d)} \) \(\large \frac{ka +k(k+1)d + a }{a * (a+kd)* (a+(k+1)d)} \) \(\large \frac{a(k+1) +k(k+1)d }{a * (a+kd)* (a+(k+1)d)} \) \(\large \frac{(k+1)(a+kd) }{a * (a+kd)* (a+(k+1)d)} \) \(\large \frac{(k+1)}{a (a+(k+1)d)} \) so, \(k+1\) th term also lies in the solution set. QED

ganeshie8 (ganeshie8):

see if that makes sense

OpenStudy (anonymous):

ok